Supposed 25 respondents were selected from a normal population with a mean of 12.5.


If the sample has a mean of 15.4 and standard deviation of 3, what percentile does the


sample mean represent?


Jon for the

1
Expert's answer
2022-04-18T15:52:42-0400

"\mu _{\bar{x}}=\mu =12.5\\\sigma _{\bar{x}}=3\\P\left( \bar{x}\leqslant 15.4 \right) =P\left( \frac{\bar{x}-\mu _{\bar{x}}}{\sigma _{\bar{x}}}\leqslant \frac{15.4-12.5}{3} \right) =\varPhi \left( 0.966667 \right) =0.833\\83.3\% percentile"